geerky42 (geerky42):

Question about limit with floor function: \[\Large \lim_{x\to\infty}\dfrac{\lfloor\sqrt{x} \rfloor^2}{x}\stackrel{?}{=}1\] I wasn't sure because there is discontinuous at every square number, but as \(x\) approaches \(\infty\), discontinuous occurs less frequent and eventually disappear. So this limit should be equal to 1. Am I right?

OpenStudy (anonymous):

Let's think about this...

OpenStudy (anonymous):

The function doesn't necessarily need to be continuous, it just needs to be continuous on a closed interval where the domain approaches.

OpenStudy (anonymous):

\[ x>\delta \implies \left|\frac{\lfloor \sqrt x\rfloor ^2}{x}-1\right|<\epsilon \]

OpenStudy (anonymous):

So I'm wondering if the drop caused by the floor function would ever push things out of our epsilon range?

geerky42 (geerky42):

Honestly, I an not familiar with limit definition and I don't see how to apply this in limit where x approaches something such as infinity.

OpenStudy (anonymous):

\[ k = \lfloor \sqrt x \rfloor \implies \sqrt{x}-1<k\leq \sqrt x \]

OpenStudy (anonymous):

We can say that \(x>0\) and so \(|x| = x\) because we can pick \(\delta >0\) if we wish.

OpenStudy (anonymous):

Hmmm, right now I'm playing around with some ideas here. Suppose: \[ \sqrt x = \lfloor \sqrt{x}\rfloor +a \]Where \(0\leq a <1\).

OpenStudy (anonymous):

Then we can plug it in \(\lfloor \sqrt x \rfloor =\sqrt x-a\). \[ \frac{(\sqrt x - a)^2}{x} = \frac{x^2-2a\sqrt x +a^2}{x} \]

OpenStudy (anonymous):

Whoops.. I messed up on that, didn't I? should be \(x\) not \(x^2\).

OpenStudy (anonymous):

\[ 1 - \frac{2a\sqrt x}{x} + \frac{a^2}{x} \]

geerky42 (geerky42):

Wow, now this is clear to me. Smart approach :)